OpenStudy (anonymous):

17. Find the length of the missing side. The triangle is not drawn to scale.

OpenStudy (anonymous):



OpenStudy (anonymous):

Do you know the formula to finding the missing sides of a triangle?

OpenStudy (anonymous):

nope lol

OpenStudy (anonymous):

i am actually really bad at math right now

OpenStudy (anonymous):

use the pythagorean formula (a² = b² + c²) a= hypotonuse b= side c= side 17² = 15² + x² 289 = 225 + x² x² = 64 Then you would Square 64 and you would get 8 :)

OpenStudy (anonymous):

So x=8 X is the missing side btw ;)

OpenStudy (anonymous):

wow ty lol

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

since u were the only one that answeres lol